# qMRI Diffusor: Quantitative T1 Mapping of the Brain using a Denoising Diffusion Probabilistic Model

### v1

@@ Abstract
Quantitative MRI (qMRI) offers significant advantages over weighted images by providing objective parameters related to tissue properties. Deep learning-based methods have demonstrated effectiveness in estimating quantitative maps from series of weighted images. In this study, we present qMRI Diffusor, a novel approach to qMRI utilising deep generative models. Specifically, we implemented denoising diffusion probabilistic models (DDPM) for T1 quantification in the brain, framing the estimation of quantitative maps as a conditional generation task. The proposed method is compared with the residual neural network (ResNet) and the recurrent inference machine (RIM) on both phantom and in vivo data. The results indicate that our method achieves improved accuracy and precision in parameter estimation, along with superior visual performance. Moreover, our method inherently incorporates stochasticity, enabling straightforward quantification of uncertainty. Hence, the proposed method holds significant promise for quantitative MR mapping.
@@ 1 Introduction
As a class of deep generative models, diffusion models have gained popularity by its impressive image generation capability. There is also a burgeoning interest in applying diffusion models within the realm of medical image analysis, encompassing various tasks such as image generation, reconstruction, registration, classification, segmentation, denoising and detection [11]. Inspired by the success of diffusion models, in this work, we present qMRI Diffusor and approach quantitative MR mapping as a generation task performed by a denoising diffusion probabilistic model (DDPM) [9] to take the advantage of diffusion models’ image generation quality. To the best of our knowledge, this is the first use of diffusion model for qMRI. In the proposed approach, the generation of quantitative maps starts with a Gaussian noise, and the results are obtained by the gradual denoising using a deep learning model. To estimate the quantitative parameters (e.g. T1) corresponding to a given series of weighted images, we set the weighted images as the condition that guides the generation process. In addition, given the inherent stochastic nature of the proposed method, uncertainty can be readily quantified by repeating the inference process [5].
@@ 2 Method
Figure 1: Illustration of qMRI Diffusor’s workflow. A U-Net model is trained to predict the noise added at a given time step. Weighted images acquired at different contrast states are set as the condition for the generation of quantitative maps. The number of channels in $y$ and $x_{t}$ should correspond to the number of weighted images used and quantitative parameters of interest respectively, and specifically here we set them as 7 and 3.
@@ 3.3 Network training
The architecture and hyperparameters of ResNet and RIM were set as the same as in [17], except that we increased RIM’s number of inference steps from 6 to 9 to boost its performance. The qMRI Diffusor was deployed based on the open-source platform MONAI Generative Models [12]. Within the qMRI Diffusor, a U-Net with time step embedding [16, 13] was configured, comprising three levels with the number of channels set to 128, 256, and 256, respectively. The original DDPM noise scheduler was used [9], which means the expected range of quantitative maps is $[-1.0,1.0]$. Hence, we scaled the training quantitative maps during pre-processing stage using $f(\boldsymbol{x})=2\times tanh(\boldsymbol{x})-1$. After training, the predicted quantitative maps were scaled back using the inverse of this function. All three methods were trained by using the same synthetic dataset with batch size of 8 and 100 epochs.
@@ 3.4 Evaluation
The phantom data and in vivo data were used to evaluate and compare ResNet, RIM and the proposed qMRI Diffusor. For qMRI Diffusor, we repeated the inference 10 times with different randomly initialised $x_{T}$ and took the mean of the outputs as the final result. Meanwhile, the voxel-wise standard deviation of the 10 outputs can act as a measurement of uncertainty. The evaluation focused on the T1 and proton density (PD) channels, as inversion efficiency can not act as a biomarker. Visual evaluation was conducted on both phantom data and in vivo data, and T1 mapping results were also quantitatively evaluated using the phantom.
@@ 4 Results
The mean and standard deviation of T1 estimations in each T1-sphere are summarised in Table 1. Within the relevant T1 range typically found in vivo (i.e., from 400 to 2000 ms), the mean value of T1 estimations using qMRI Diffusor has the closest match to the ground truth values. Additionally, these estimations exhibit the smallest standard deviation in the ROIs. For T1 lower than 500 ms, qMRI Diffusor exhibits a bias towards values larger than the ground truth, with a standard deviation mostly smaller than that of RIM but larger than that of ResNet.
@@
GT ResNet RIM Diffusor GT ResNet RIM Diffusor
@@
Fig. 3 shows examples of T1 and PD estimations for in vivo data. In the T1 estimation, qMRI Diffusor shows the ability to preserve small anatomical structures, while blurriness around small structures are observed from both ResNet’s and RIM’s results. Within the brain mask, high uncertainty values given by qMRI Diffusor are observed in the high T1 region (above 2000 ms), while the uncertainty values in gray matter/white matter region is approximately homogeneous. The PD estimations from all methods are visually similar without outliers or noticeable artifacts, while RIM’s estimation is lower in the white matter and qMRI Diffusor’s estimation is slightly higher overall.
@@ 5 Discussion
According to the evaluation results, the proposed method trained purely on synthetic data generalises well to data from real scans, and shows resilience to artifacts in the input weighted images. The test with phantom shows that qMRI Diffusor achieves the best visual results for T1 mapping by providing clean edges of spheres and uniform T1 values within the spheres, as well as higher accuracy and precision for T1 values in the range typically found in vivo. Though qMRI Diffusor’s performance for T1 values less than 500 ms is admittedly suboptimal, this could be attributed to the fact that such T1 values are rare in the synthetic human brain training dataset, and the fact that qMRI Diffusor is not physics model-based, which may limit the ability to extrapolate outside of the range of values seen during training. However, it could be improved by using a training dataset with a larger T1 range. The proposed method also provides superior visual performance on the in vivo data compared to ResNet and RIM, and the estimated T1 maps using qMRI Diffusor accurately depicted anatomical structures including the fine details.
One significant advantage of the proposed method is that it readily allows uncertainty quantification, which adds trustworthiness and interpretability to this method. For other methods, uncertainty quantification is either infeasible or dependent on separate operations (e.g. by adding dropout layers in the neural network [8]). The uncertainty quantification by using voxel-wise standard deviation of repeated inferences is meaningful, as the result reconstructed from a randomly initialised $x_{t}$ can be viewed as a sample from the posterior distribution, given the training data the model has seen. It is noticeable that high-uncertainty region given by qMRI Diffusor overlaps with its error-prone areas (e.g. edge of spheres, high T1 regions), indicating the meaningfulness and usefulness of the uncertainty information provided by our method.
@@ 6 Conclusion
We propose a novel quantitative MR mapping method, qMRI Diffusor, which frames the mapping process as a conditional generation task using a denoising diffusion probabilistic model (DDPM). The visual and quantitative evaluation on real scans demonstrates the superior performance of the proposed method, highlighting its potential in clinical use. Particularly, the proposed method allows readily uncertainty quantification. In conclusion, the proposed method could be a promising tool for qMRI.

### v2

# qMRI Diffuser: Quantitative T1 Mapping of the Brain using a Denoising Diffusion Probabilistic Model

@@ Abstract
Quantitative MRI (qMRI) offers significant advantages over weighted images by providing objective parameters related to tissue properties. Deep learning-based methods have demonstrated effectiveness in estimating quantitative maps from series of weighted images. In this study, we present qMRI Diffuser, a novel approach to qMRI utilising deep generative models. Specifically, we implemented denoising diffusion probabilistic models (DDPM) for T1 quantification in the brain, framing the estimation of quantitative maps as a conditional generation task. The proposed method is compared with the residual neural network (ResNet) and the recurrent inference machine (RIM) on both phantom and in vivo data. The results indicate that our method achieves improved accuracy and precision in parameter estimation, along with superior visual performance. Moreover, our method inherently incorporates stochasticity, enabling straightforward quantification of uncertainty. Hence, the proposed method holds significant promise for quantitative MR mapping.
@@ 1 Introduction
As a class of deep generative models, diffusion models have gained popularity by its impressive image generation capability. There is also a burgeoning interest in applying diffusion models within the realm of medical image analysis, encompassing various tasks such as image generation, reconstruction, registration, classification, segmentation, denoising and detection [11]. Inspired by the success of diffusion models, in this work, we present qMRI Diffuser and approach quantitative MR mapping as a generation task performed by a denoising diffusion probabilistic model (DDPM) [9] to take the advantage of diffusion models’ image generation quality. To the best of our knowledge, this is the first use of diffusion model for qMRI. In the proposed approach, the generation of quantitative maps starts with a Gaussian noise, and the results are obtained by the gradual denoising using a deep learning model. To estimate the quantitative parameters (e.g. T1) corresponding to a given series of weighted images, we set the weighted images as the condition that guides the generation process. In addition, given the inherent stochastic nature of the proposed method, uncertainty can be readily quantified by repeating the inference process [5].
@@ 2 Method
Figure 1: Illustration of qMRI Diffuser’s workflow. A U-Net model is trained to predict the noise added at a given time step. Weighted images acquired at different contrast states are set as the condition for the generation of quantitative maps. The number of channels in $y$ and $x_{t}$ should correspond to the number of weighted images used and quantitative parameters of interest respectively, and specifically here we set them as 7 and 3.
@@ 3.3 Network training
The architecture and hyperparameters of ResNet and RIM were set as the same as in [17], except that we increased RIM’s number of inference steps from 6 to 9 to boost its performance. The qMRI Diffuser was deployed based on the open-source platform MONAI Generative Models [12]. Within the qMRI Diffuser, a U-Net with time step embedding [16, 13] was configured, comprising three levels with the number of channels set to 128, 256, and 256, respectively. The original DDPM noise scheduler was used [9], which means the expected range of quantitative maps is $[-1.0,1.0]$. Hence, we scaled the training quantitative maps during pre-processing stage using $f(\boldsymbol{x})=2\times tanh(\boldsymbol{x})-1$. After training, the predicted quantitative maps were scaled back using the inverse of this function. All three methods were trained by using the same synthetic dataset with batch size of 8 and 100 epochs.
@@ 3.4 Evaluation
The phantom data and in vivo data were used to evaluate and compare ResNet, RIM and the proposed qMRI Diffuser. For qMRI Diffuser, we repeated the inference 10 times with different randomly initialised $x_{T}$ and took the mean of the outputs as the final result. Meanwhile, the voxel-wise standard deviation of the 10 outputs can act as a measurement of uncertainty. The evaluation focused on the T1 and proton density (PD) channels, as inversion efficiency can not act as a biomarker. Visual evaluation was conducted on both phantom data and in vivo data, and T1 mapping results were also quantitatively evaluated using the phantom.
@@ 4 Results
The mean and standard deviation of T1 estimations in each T1-sphere are summarised in Table 1. Within the relevant T1 range typically found in vivo (i.e., from 400 to 2000 ms), the mean value of T1 estimations using qMRI Diffuser has the closest match to the ground truth values. Additionally, these estimations exhibit the smallest standard deviation in the ROIs. For T1 lower than 500 ms, qMRI Diffuser exhibits a bias towards values larger than the ground truth, with a standard deviation mostly smaller than that of RIM but larger than that of ResNet.
@@
GT ResNet RIM Diffuser GT ResNet RIM Diffuser
@@
Fig. 3 shows examples of T1 and PD estimations for in vivo data. In the T1 estimation, qMRI Diffuser shows the ability to preserve small anatomical structures, while blurriness around small structures are observed from both ResNet’s and RIM’s results. Within the brain mask, high uncertainty values given by qMRI Diffuser are observed in the high T1 region (above 2000 ms), while the uncertainty values in gray matter/white matter region is approximately homogeneous. The PD estimations from all methods are visually similar without outliers or noticeable artifacts, while RIM’s estimation is lower in the white matter and qMRI Diffuser’s estimation is slightly higher overall.
@@ 5 Discussion
According to the evaluation results, the proposed method trained purely on synthetic data generalises well to data from real scans, and shows resilience to artifacts in the input weighted images. The test with phantom shows that qMRI Diffuser achieves the best visual results for T1 mapping by providing clean edges of spheres and uniform T1 values within the spheres, as well as higher accuracy and precision for T1 values in the range typically found in vivo. Though qMRI Diffuser’s performance for T1 values less than 500 ms is admittedly suboptimal, this could be attributed to the fact that such T1 values are rare in the synthetic human brain training dataset, and the fact that qMRI Diffuser is not physics model-based, which may limit the ability to extrapolate outside of the range of values seen during training. However, it could be improved by using a training dataset with a larger T1 range. The proposed method also provides superior visual performance on the in vivo data compared to ResNet and RIM, and the estimated T1 maps using qMRI Diffuser accurately depicted anatomical structures including the fine details.
One significant advantage of the proposed method is that it readily allows uncertainty quantification, which adds trustworthiness and interpretability to this method. For other methods, uncertainty quantification is either infeasible or dependent on separate operations (e.g. by adding dropout layers in the neural network [8]). The uncertainty quantification by using voxel-wise standard deviation of repeated inferences is meaningful, as the result reconstructed from a randomly initialised $x_{t}$ can be viewed as a sample from the posterior distribution, given the training data the model has seen. It is noticeable that high-uncertainty region given by qMRI Diffuser overlaps with its error-prone areas (e.g. edge of spheres, high T1 regions), indicating the meaningfulness and usefulness of the uncertainty information provided by our method.
@@ 6 Conclusion
We propose a novel quantitative MR mapping method, qMRI Diffuser, which frames the mapping process as a conditional generation task using a denoising diffusion probabilistic model (DDPM). The visual and quantitative evaluation on real scans demonstrates the superior performance of the proposed method, highlighting its potential in clinical use. Particularly, the proposed method allows readily uncertainty quantification. In conclusion, the proposed method could be a promising tool for qMRI.